calculate the critical angle and an angle of entry for an optical fibre having core of refractive index 1.50 and cladding of refractive index 1.48a. 30 degree b. 14.2 degreec. 5 deg d. 45 degree

Grade:12th pass

1 Answers

Khimraj
5 years ago
Let n1 = refractive index of core = 1.50
n2 = refractive index of cladding = 1.48
critical angle is given by
sin\Theta _{c} = n2/n1
So \Theta _{c} = sin-1(1.48/1.50) = 80.63o
and angle of entry is given by
sin\alpha = \sqrt{n_{1}^{2} - n_{2}^{2}}
\alpha = 14.2o